OpenStudy (anonymous):

MEDAL The sum of a number and six is multiplied by negative four. The result is eight. What is the number?

OpenStudy (anonymous):

well first you set up an equation

OpenStudy (anonymous):

\[(x+6)*-4\]

OpenStudy (anonymous):

=8

OpenStudy (anonymous):

ok so thats what u have

OpenStudy (anonymous):

so divide both side by negative 4

OpenStudy (anonymous):

\[\frac{ (x+6)*-4 }{ -4 }=\frac{ 8 }{ -4 }\]

OpenStudy (anonymous):

can you tell me the answer to that?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

so you have now (x+6)=-2

OpenStudy (anonymous):

so now you do -2+6

OpenStudy (anonymous):

and that should tell you what x is